OpenStudy (anonymous):

The square root of y squared = ? a. y b. y2 c. y3 d. y4

OpenStudy (solomonzelman):

\[\sqrt{y^2}=?\]

OpenStudy (solomonzelman):

Do you know?

OpenStudy (anonymous):

no i dont know

OpenStudy (solomonzelman):

\[\sqrt[a]{b^a}=b\]

OpenStudy (anonymous):

The square root of a number squared is equal to the absolute value of that number.